Aggregator

Autolog save

21 hours 38 minutes ago
Log Number: 4468896Logbook(s): Tags: Body: saver ran from opsl05

Saved: lrichard 02-06-26 15:54:36 {A:5 B:3 C:1 E:340 A=BS B=BS C=BS D=BS 0R08=0.00 0R08 dump in. Start of swing; After Max completed his work} { } ALL

Comments:
lrichard

CREW CHIEF SIGN IN

21 hours 40 minutes ago
Log Number: 4468895Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 02-06-26
Time : 15:38:44
Day : Friday
Shift : Swing

Crew Chief : Richardson
Program Deputy: Mock
Team : X
Arm : RADCON
Designated SSO: Schaumloffel
LERF Operator:
Operators : Schaumloffel

This stamp certifies that the above Crew Chief has read/reviewed/received the following:
lrichard

BSList #117047: measurement to optimize HWien focusing works

21 hours 57 minutes ago
Log Number: 4468893Logbook(s): Entry Makers: brukerBody: 

Retracted A1. PB can stay on, the comparison should still work.
VWien off for now, MFX2I01 = 208, MFX1I03 = 196.
MFX0I01 = 0.
HWien is still limited to 60 degrees pending DRVH adjustment, so "on" means 60 degrees today.

Measuring beam size on 0I02 viewer for HWien on/off as a function of MFG1I04 setpoints under the constraint that B = -A, avoiding coupling for now.

bruker

iocsoftw3spce restored

22 hours 1 minute ago
Log Number: 4468888Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-02-06 15:27:00 The output from the myRestore execution was as follows: iocsoftw3spce systemEnum@2026-02-06 15:27 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -i iocsoftw3spce (success) PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 Procedure return status: Success
alarms

APPROVED ONGOING BSY ion pump swap

22 hours 5 minutes ago
Log Number: 4468886Logbook(s): Body: 

Spoke with Greg Marble.

Greg, with additional SRFVAC support, will begin Monday morning swapping five ion pumps: VIP2C00A, VIPAT09, and VIP3C{01,03,03A}. VIP3C04 is elevated but still in the "green" range. Swapping the upstream pumps will bring it down even further. VIPAT09 is railed even after the earlier nT-VIP swaps, and with its closeness to the others makes good sense to address together with the 2C and 3C ones.

Anticipates this job taking 2-3 days. This works fine, even if it spills over another day or two.

moser

Autolog save

22 hours 12 minutes ago
Log Number: 4468885Logbook(s): Tags: Body: saver ran from opsl06

Saved: bruker 02-06-26 15:20:28 {A:5 B:3 C:1 E:340 A=Tune B=BS C=BS D=BS 0R08=0.00 0R08 dump in.
bruker

HOLD BSY ion pump swap

22 hours 47 minutes ago
Log Number: 4468877Logbook(s): Body: 

I appreciate the initiative but need more information before you can proceed.
* The original list of ion pumps was VIP3C {01,03,03A, and 04} -- why are we including AT09 now? The stretch goal was VIP2C00A, is that no longer in plan?
* Once you begin, how long do you anticipate having cleanrooms up? There is a VIP tour coming through Monday 2/16, with a dry run occurring Thursday 2/12. Would this interfere with the vacuum work that has to be done?

moser

2L26-4 Solenoid Low Current

22 hours 54 minutes ago
Log Number: 4468876Logbook(s): Entry Makers: kernsBody: 

Replaced the klystron Sorenson PS Tuesday after it was over-temping and shutting off. Came back on after replacing it with 15.6a, instead of ~ 31.3a. Might need to replace it again for this current readback problem that the last unit didn't have. Rechecked all cable connections and didn't see a problem.

kerns

iocsoftmdaq1 restored

23 hours ago
Log Number: 4468875Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2026-02-06 14:28:33 The output from the myRestore execution was as follows: iocsoftmdaq1 systemEnum@2026-02-06 14:28 2(2) restored rtpcellxport@2026-02-06 14:28 24(24) restored rtpcellCalc@2026-02-06 14:28 8(8) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -i iocsoftmdaq1 (success) PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 PVs modified: 0 Procedure return status: Success
alarms

VIP2L258 HV Cable Needs Repair

23 hours 23 minutes ago
Log Number: 4468865Logbook(s): Body: 

Before 2L25 cryomodule was installed I load tested the HV cables for each pump. VIP2L258 showed to be overloading with nothing connected to it. This is indicative of rad damaged cable and will need to be re-terminated. An ATLis and epas will be made for the week to be done next week.

mathewam

HYST AREA 7-0 Cycles Magnets

23 hours 36 minutes ago
Log Number: 4468862Logbook(s): Body: 

HYST AREA Version 7-0 has successfully cycled the following:

MQB1L23
MQB1L24
MQB1L25
MQB1L26
MQB1L27
MQB1L28
MQB1S01
MBT1S01H
MBT1S01V
MQA3S01
MBC3S01H
MBC3S01V

Total Time: 1.550 mins

diggs

Shift Information

Shift Day
CEBAF Program
MCC Crew Chief Blackshaw
MCC Operators Thomas
Program Deputy Mock
Announcements

Access Status

Recent PSS History

recent pss history graph